OpenStudy (anonymous):

Solve for a. y=ax^2+bx+c Please help!

jimthompson5910 (jim_thompson5910):

Subtract c from both sides Then subtract bx from both sides

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (lgbasallote):

a celebratory party?

jimthompson5910 (jim_thompson5910):

...well before that lol

OpenStudy (anonymous):

divide x^2?

jimthompson5910 (jim_thompson5910):

yes, divide both sides by x^2

OpenStudy (anonymous):

I'm a little confused about how to put it all together

jimthompson5910 (jim_thompson5910):

\[\Large y=ax^2+bx+c\] \[\Large y-c=ax^2+bx\] \[\Large y-c-bx=ax^2\] \[\Large \frac{y-c-bx}{x^2}=a\] \[\Large a=\frac{y-c-bx}{x^2}\]

OpenStudy (anonymous):

a=c-y/x^2-bx?

jimthompson5910 (jim_thompson5910):

close, but see above

OpenStudy (anonymous):

is that the same as a=-b/x+y/x^2-c/x^2

jimthompson5910 (jim_thompson5910):

yes that's another way to state the answer

OpenStudy (anonymous):

Thank you so much!!!

jimthompson5910 (jim_thompson5910):

np